OpenStudy (anonymous):

How many different 12-member juries can be made from 150 people? Please explain how. Thank you (:

OpenStudy (anonymous):

150 choose 12 or 150!/(12!(138!))

OpenStudy (nubeer):

try 150C12 means combination of 12 from 150

OpenStudy (anonymous):

it is a huge number you are asked to compute \(\dbinom{150}{12}\) read "150 choose 12" the number of ways you can pick 12 out of 150 you can try to compute \[\dbinom{150}{2}=\frac{150!}{12!138!}\] if you like, but i would do this http://www.wolframalpha.com/input/?i=150+choose+12

OpenStudy (anonymous):

the computation would be \[\frac{150\times 149\times 148\times 147\times 146\times 145\times 144\times 143\times 1141\times 140\times 139}{12\times 11\times 10\times 9\times 8\times 7\times 6\times 5\times 4\times 3\times 2 }\]

OpenStudy (anonymous):

Thank you so much (: